Operations Report – February 20th

Crew 222 Operations Report 20Feb2020

SOL: 4

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Wednesday generator:

Off at 6:30am

On at 9:00pm SOC was 83%

Thursday generator:

Off at 5:30am

On at 9pm SOC was XX%

Diesel Reading – 75%

Station Propane Reading – 34% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149356.82 units

Water (static tank) – 387 gallons

Static to Loft Pump used – yes

Water in GreenHab – 108.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 144.3

Beginning charge: 100%

Ending charge: 94%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: no

Hours: 132.0

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes

Opportunity rover used: no

Hours: 69.9

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes

Curiosity rover used: yes

Hours: 138.2

Beginning charge: 100%

Ending charge: 69%

Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes

Reason for use: Run during the engineer EVA for the battery.

Oil Added? no

ATV Fuel Used: N/A

Hours the ATVs were Used today: 10 minutes

Notes on ATVs: The second one from the Hab doesn’t have anymore battery.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com. All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the EVA. The spacesuit number 10 has stopped ventilating during the EVA this morning but the lights were still on. I tried to fix this during this afternoon and I’ll see if it works tomorrow.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site